OpenStudy (anonymous):

If f(x) = x2 + 7, what is the value of f(3)?

OpenStudy (texaschic101):

f(x) = x^2 + 7 f(3) = 3^2 + 7 f(3) = 9 + 7 f(3) = 16 all you do is sub in 3 for x and solve